A black spherical ball (epsilon = 0.95) at 125°C is suspended in air. It loses heat by convection to the surrounding air at 25°C with an associated convection heat transfer coefficient of 12 W·m-2·K-1 and by radiation to the surrounding surfaces at 15°C. Assuming steady state conditions, determine the heat transfer flux due to convection and radiation.
Answer #1

Answer & Ciren E=0.95 0-5.staw w T= 125c-3a8k. To = 150-2856 Heat loses by radiation - 5 As To - To & Heat toeste, then cleueheat tawater duto radiation and Convention Q = Qr the = 980.99548200 À = 2180.995 W/m2